OpenStudy (anonymous):

hello i would like to know if you guys can help me in some basic equivalent fractions. im trying to help my son.

OpenStudy (saifoo.khan):

Hi, Sure post your problems!

OpenStudy (anonymous):

_/4=24/32

OpenStudy (saifoo.khan):

Let's suppose that number be "x" \[\frac{x}{4} = \frac{24}{32}\] Now what we have to do is get rid of the "4" on the left side.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

To do that we have to multiply "4" on both sides, \[4 \times \frac{x}{4} = \frac{24}{32} \times 4\]\[\cancel 4 \times \frac{x}{\cancel 4} = \frac{24}{32} \times 4\]

OpenStudy (saifoo.khan):

So now we are left with, \[x = \frac{24}{32} \times 4\] Can you solve now?

OpenStudy (anonymous):

im lost x has to be a number that is multipy equal

OpenStudy (saifoo.khan):

Actually what we did is we supposed the blank as x.

OpenStudy (saifoo.khan):

So we can find that number by making "x" the subject and then you will get the answer.

OpenStudy (anonymous):

ok give me the give me the number so i can see how its done

OpenStudy (saifoo.khan):

Join me here, i will tell you step by step: http://www.twiddla.com/763478